OpenStudy (anonymous):

solve the inequality (6x+21)(94x-63)<=0 write answer in interval notation

OpenStudy (anonymous):

i got \[\frac{ -7 }{ 2 }\ge x \ge \frac{ 63 }{ 94 }\]

OpenStudy (anonymous):

then in notation i wrote (-infinity,-7/2]U[63/94,infinity)

OpenStudy (anonymous):

and it says its wrong :/

OpenStudy (solomonzelman):

ohh, you put the wrong interval, but your solution was good.

OpenStudy (solomonzelman):

the interval should be: \(\large\color{black}{ [-\frac{\LARGE 7}{\LARGE 2} ,\frac{\LARGE 63}{\LARGE 94} ~ ] }\)

OpenStudy (solomonzelman):

the reason is, because your x solution is from (and including,) \(\large\color{black}{ -\frac{\LARGE 7}{\LARGE 2} }\) and till (and including,) \(\large\color{black}{ \frac{\LARGE 63}{\LARGE 94} ~. }\)

OpenStudy (solomonzelman):

what you wrote for interval notation would satisfy the: \(\large\color{black}{(6x+21)(94x-63)\ge0 }\)